"x+\sqrt{12-\sqrt{x}}=12"

Solve this equation for x.

Please provide the full steps.


1
Expert's answer
2022-05-01T17:01:21-0400
"0\le x\le12"

"\sqrt{12-\sqrt{x}}=12-x""\bigg(\sqrt{12-\sqrt{x}}\bigg)^2=(12-x)^2"

"12-\sqrt{x}=144-24x+x^2"

"x^2-24x+\sqrt{x}+132=0"

"(\sqrt{x})^4-24(\sqrt{x})^2+\sqrt{x}+132=0"

"(\sqrt{x}+4)(\sqrt{x}-3)((\sqrt{x})^2-\sqrt{x}-11)=0"

"\sqrt{x}+4>0"

"\sqrt{x}-3=0=>x=9"

"(\sqrt{x})^2-\sqrt{x}-11=0"

"D=(-1)^2-4(1)(-11)=45"

"\sqrt{x}=\dfrac{1\pm\sqrt{45}}{2}"

Since "\sqrt{x}\ge0," we take


"\sqrt{x}=\dfrac{1+\sqrt{45}}{2}"

"x=\dfrac{(1+3\sqrt{5})^2}{4}"

"x=\dfrac{23+3\sqrt{5}}{2}>12, does\ not\ satisfy"

"\{9\}"